User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Jay

Next meeting chair: Eloise

Attended: Mas, Katie, Eloise, Earl, Prad, Jamie, Patrick, Lim

Apologies: Simin, Tiege,

General Issues

  • CAL controller failure for HO 26 during AOV051
    • one of amplifier stage stopped working. it's been fixed now. it's been repaired now. need to confirm whether it's working fine or not
  • Autocorrelations have been funky for couple of H0 26 experiments ( AOV051 & R1966)
    • might have been something to do with proc files. it should be fixed now. please keep an eye if any problem occurs in future

Handover Issues

  • Hb 12:
  • AOV051: Experiment started with recording in only one stream. Check and make sure that it's record each scan in three stream(sxy,xx,xy) at the start of experiment.
    • Jamie and Guifre working on smoothing the whole procedure of running different scripts for mixed mode exp. but till then keep an eye on issues and upadte the handover/setup notes for mixed modes whenever you find something useful.
  • H0 26
  • Mk5 wasn't communicating with fs. restarting drive dv seems to fixed it but still no recording. Maybe module issue? Power cycle of dbbc done the job at last.
/home/www/auscope/opswiki/data/pages/operations/agenda2020_09_24.txt · Last modified: 2020/10/08 02:22 by Jayender Kumar